PPP - Purchasing Power Parity

Suppose that the current Swiss franc to U.S. dollar spot exchange rate is $:SFr = 1.60 (i.e., 1.60 SFr per U.S. dollar or 1.60 SFr/$). The expected inflation over the coming year is 2% in Switzerland and 5% in the US. According to the purchasing power parity, what is the expected value of the Swiss franc to U.S. dollar spot exchange rate a year from now?
Answer #1

AS per PPP theory

\large Forward Rate = Spot Rate * (\frac{1+Swiss Inflation Rate}{1+ US Inflation Rate })

\large Forward Rate = 1.6 * (\frac{1+0.02}{1+ 0.05 })

= 1.5543

Forward Rate is 1 USD = 1.5543 SFr

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.